Find S(n) for the arithmetic series in which a(1)=-11, a(n)=13, and n=7.
Thank you

Found 2 solutions by venugopalramana, fractalier:
Answer by venugopalramana(3286)

Find S(n) for the arithmetic series in which a(1)=-11, a(n)=13, and n=7.
Thank you
SN=(N/2){A1+AN}=(7/2)(-11+13)=7

Answer by fractalier(2101)

The formula for the sum of an arithmetic series is
S = (n/2)(a(1) + a(n))
S = (7/2)(-11 + 13)
S = (7/2)(2)
S = 7